OpenStudy (anonymous):

find the slop of the line that passes through the giving points. (-13,7) and (17,-11)

OpenStudy (aroub):

-11-7/17+13 =-17/30

OpenStudy (amistre64):

slope is the change in y values over the change in x values \[\frac{\triangle y}{\triangle x}\]

OpenStudy (amistre64):

subtract your points and stack y/x (-13, 7) -( 17,-11) ---------- -30, 18; slope = y/x = -18/30 = -9/15